If the square of 3 more than a number is 9, find the numbers?
Answer by bucky(2189)

Let x be the unknown number.
.
Then 3 more than this unknown number is x + 3
.
And squaring this quantity is:
.
(x + 3)^2
.
The problem tells you that this is to equal 9. In equation form this is:
.
(x + 3)^2 = 9
.
Squaring out the left side results in this equation becoming:
.
x^2 + 6 x + 9 = 9
.
Subtract 9 from both sides results in:
.
x^2 + 6x = 0
.
Factor the left side:
.
x(x + 6) = 0
.
Note that the left side of the equation will be zero if either of the two factors is
equal to zero ... because multiplying by a zero factor will result in an answer of zero.
.
So for the two answers, either:
.
x = 0
.
or
.
x + 6 = 0
.
In this second case, subtracting 6 from both sides results in:
.
x = -6
.
So the two answers are x = 0 and x = -6.
.
To check these two answers, return to the equation:
.
(x + 3)^2 = 9
.
Try the first answer we got ... x = 0. Substituting this value of x into the equation
and you get:
.
(0 + 3)^2 = 9
.
This simplifies to 3^2 = 9 and that, of course, checks out.
.
Next in the equation substitute -6 for x and you get:
.
(-6 + 3)^2 = 9
.
Combining the two numbers in the parentheses leads to:
.
(-3)^2 = 9
.
And this too is correct because squaring -3 results in +9.
.
So our answers of x = 0 and x = -6 are correct.
.
Hope this helps you to understand the problem and how to work out an answer.
.